Run the following OPTP command:
Optp cmd : 0813
01 04 00 e7 00 bf
Total records :1
In the foregoing commands, “1c” means the slot is 28, the returned value “00e7” indicates the upper threshold, and “00bf” indicates the lower threshold. We find that the set threshold is 19.1~23.1°C. However, the actual performance value is 23.7°C. It indicates that the alarm reporting is normal, and the threshold value of the optical module is set improperly. As confirmed, the site cannot load the threshold value of modifying the optical module. The TEM_HA alarm can be shielded to solve the problem temporarily. To solve the problem thoroughly, the board needs to be replaced.